HIGH RESOLUTION PROBES OF COUPLED STATES: ZEEMAN QUANTUM BEAT STUDIES OF OD(A^{2}Σ^{+},v=0,1,2,3) AND OH(A^{2}Σ^{+},v=0,1,2)

J. XIN, I. IONESCU, D. KUFFEL, AND \underline{S. A. REID}, Marquette University, Dept. of Chemistry, Milwaukee, WI USA 53233.

We will report on Zeeman quantum beat studies of rovibronic levels in the A^{2}Σ^{+} state of the hydroxyl and hydroxyl-d radicals. For each system the data extends to the onset of predissociation induced via spin-orbit coupling with a repulsive 4Σ^{−} state, which occurs for OH at v=2 and for OD at v=3. Our study focused on identifying direct spectroscopic signatures of this coupling as revealed in the weak field Lande g-factors. For OD, we were also able to determine the excited state hyperfine parameters, which are a sensitive probe of the excited state electronic structure, and trends in these parameters will be discussed.